The International Air Transport Association surveys business travelers to develop quality ratings for transatlantic gateway airports. The maximum possible rating is 10. Suppose a simple random sample of 50 business travelers is selected and each traveler is asked to provide a rating for the Miami International Airport. The ratings obtained from the sample of 50 business travelers follow. COL A-Z oooooooo The International Air Transport Association surveys business travelers to develop quality ratings for transa

Teatures of Excel have been disabled because ith X fx Rating non con Data Sheet2 Sheet3 Ready
Answer #1
Values ( X ) Σ ( Xi- X̅ )2
6 0.1156
4 5.4756
6 0.1156
8 2.7556
7 0.4356
7 0.4356
6 0.1156
3 11.1556
3 11.1556
8 2.7556
10 13.3956
4 5.4756
8 2.7556
7 0.4356
8 2.7556
7 0.4356
5 1.7956
9 7.0756
5 1.7956
8 2.7556
4 5.4756
3 11.1556
8 2.7556
5 1.7956
5 1.7956
4 5.4756
4 5.4756
4 5.4756
8 2.7556
4 5.4756
5 1.7956
6 0.1156
2 18.8356
5 1.7956
9 7.0756
9 7.0756
8 2.7556
4 5.4756
8 2.7556
9 7.0756
9 7.0756
5 1.7956
9 7.0756
7 0.4356
8 2.7556
3 11.1556
10 13.3956
8 2.7556
9 7.0756
6 0.1156
Total 317 229.22

Mean X̅ = Σ Xi / n
X̅ = 317 / 50 = 6.34
Sample Standard deviation SX = √ ( (Xi - X̅ )2 / n - 1 )
SX = √ ( 229.22 / 50 -1 ) = 2.1629

Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.05 /2, 50- 1 ) = 2.01
6.34 ± t(0.05/2, 50 -1) * 2.1629/√(50)
Lower Limit = 6.34 - t(0.05/2, 50 -1) 2.1629/√(50)
Lower Limit = 5.73
Upper Limit = 6.34 + t(0.05/2, 50 -1) 2.1629/√(50)
Upper Limit = 6.95
95% Confidence interval is ( 5.723 , 6.95 )